OpenStudy (anonymous):

what is the slope of x+3y=6 I think it is -1/3 but i just want to make sure

OpenStudy (anonymous):

Solve x+3y=6 for y, what do you get?

OpenStudy (ryan123345):

put -x so the problem becomes 3y=-x + 6 then to get rid of the 3y you divide 3 on each side you get y = -x/3 + 6/3 so the answer is 7 = -x/3 + 2